The Temperature Dependence of Chemical Equilibrium

For an equilibrium reaction:

A + B = C + D


dynamically generated plot

Given the above relationship, use the graph to identify which of the following statements are "true" or "false" in this specific case.

All statements refer to the forward reaction (proceeding left to right, as given) corresponding to the equilibrium constant K.

1)The reaction is spontaneous at the highest temperatures but not at lower temperatures.
2)deltaS° for the reaction is negative.
3)deltaG° is negative at low temperatures but becomes positive at higher temperatures.
4)The reaction is exothermic.
5)As the temperature is decreased, the equilibrium position shifts in favour of the products.
Answer #1

1)The reaction is spontaneous at the highest temperatures but not at lower temperatures.

as 1/T increases, T decreases.. so... the graph shows that as T decreases ln(K) decreases so....

At high T, expect K >> 1; so this is TRUE

2)deltaS° for the reaction is negative.

From the plot:

y-intercept is positive, so this means

positive = dS/R

so dS must be positive, the argument is FALSE

Q3.

dG = dG°-RT*ln(K)

dG°, by definition, will not change as we change T, since dG° is at standard conditions, the argument is FALSE

Q4.

slope = - dH/R

since slope is negative then

-X = -dH/R

meaning

X*R = dH

so the dH is positive

meaning it is ENDOTHERMIC, the arguemtn is false

Q5

As T decreases... ln(K) increases, meaning it favours products... this is TRUE